A condensed balance sheet for Bradford Corporation prepared at the end of the year appears as follows. Assets Liabilities & Stockholders' Equity Cash $ 114,000 Notes payable (due in 6 months) $ 48,000 Accounts receivable 100,000 Accounts payable 97,000 Inventory 260,000 Long-term liabilities 316,000 Prepaid expenses 60,000 Capital stock, $5 par 300,000 Plant & equipment (net) 570,000 Retained earnings 433,000 Other assets 90,000 Total $ 1,194,000 Total $ 1,194,000 During the year, the company earned a gross profit of $1,116,000 on sales of $2,950,000. Accounts receivable, inventory, and plant assets remained almost constant in amount throughout the year, so year-end figures may be used rather than averages.

a. Compute the current ratio. (Round your answer to 2 decimal place.)

b. Compute the quick ratio. (Round your answer to 2 decimal place.)

c. Compute the working capital.

d. Compute the debt ratio. (Round your percentage answers to nearest whole percent. i.e. 0.1234 as 12%.)

e. Compute the accounts receivable turnover (all sales were on credit). (Round your answer to 2 decimal places.)

f. Compute the inventory turnover. (Round your answer to 2 decimal places.)

g. Compute the book value per share of capital stock. (Round your answer to 2 decimal places.)
Answer #1
Requirement a.
Current Ratio = Current assets / Current liabilities
= (114000+100000+260000+60000)/ (48000+97000)
= 534000/ 145000
= 3.68
Requirement b.
Quick ratio = Quick assets / Current liabilities
= (114000+100000)/ (48000+97000)
= 214000/ 145000
= 1.48
Requirement c.
Working capital = Current assets - Current liabilities
= (114000+100000+260000+60000) - (48000+97000)
= 534000 - 145000
= 389,000
Requirement d.
Debt ratio = Total debt / Total Assets
= (48000+97000+316000)/1194000
= 39%
Requirement e.
Accounts receivable turnover = Annual net credit sales / Average accounts receivable
=2,950,000/100,000
=29.50
Requirement f.
Inventory turnover = Cost of goods sold / Average inventory
= (2,950,000-1,116,000)/260,000
= 7.05